OpenStudy (anonymous):

Solve: 5x - 2x + 6 = 18 a. x = -4 b. x = 4 c. x = 6 d. x = 8

hero (hero):

3x+6 = 18 3x = 12 x = 4

OpenStudy (anonymous):

b. x=4

OpenStudy (anonymous):

If you simplify the like terms i.e. terms with 'x' here you will get 3x Hence the question becomes 5x - 2x + 6 = 18 3x +6 = 18 we need to isolate the x; So take away 6 from both sides 3x =12 To find 'x', divide both sides by 3 we get x = 4